OpenStudy (friendzone):

W = pt + 105 A wrestler is trying to gain weight for an upcoming match. His current weight is 105. In the formula shown, W = weight, p = pounds and t = time(weeks). He has 6 weeks before his match and has a goal weight of 123 pounds. Determine the minimum number of pounds he must gain per week to meet his goal. A) 1 pound B) 2 pounds C) 3 pounds D) 4 pounds

OpenStudy (anonymous):

Plug them in for 'p' Here, 't' is 6, because we're looking at the time for 6 weeks. W = (1)(6) + 105 Multiply: W = 6 + 105 Add: W = 111 This looks too small, try another number the same way. @friendzone

OpenStudy (texaschic101):

try plugging in 3 for p and see what you get

OpenStudy (anonymous):

^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^

OpenStudy (anonymous):

W = (3)(6) + 105 What's 3 * 6? @friendzone

OpenStudy (friendzone):

got 3